# A Wallace semigroup whose every finite power is countably compact

### 01 Mar 2024 v1

@@ 1.2. Basic results and notation
A free ultrafilter $p$ is said to be selective (or Ramsey) if for every $f:[\omega]^{2}\rightarrow 2$ there exists $A\in p$ such that $f|_{[A]^{2}}$ is constant. It is well known that selective ultrafilters are exactly the minimal ultrafilter in the Rudin-Keisler order, which is the pre-order defined in the collection of all free ultrafilters defined by $p\leq_{\operatorname{RK}}q$ if there exists a function $f:\omega\to\omega$ such that $p=\{f^{-1}[A]:A\in q\}$. Two selective ultrafilters are said to be incomparable if they are incomparable with respect to $\leq_{\operatorname{RK}}$.
@@ 4. $\mathbb{H}$ has no semigroup topology which makes its $\omega$-th power countably compact
We prove something more general. For that purpose, we say that an Archimedian semigroup with identity is a $4$-uple $(T,+,0,\leq)$ such that $(T,+,0)$ is an Abelian semigroup with identity 00, $\leq$ is a total order such that for all $a,b,c,d\in A$, if $a\leq b$ and $c\leq d$, then $a+c\leq b+d$, and for all $a>0$ and $b\in T$, there exists $n\in\mathbb{N}$ such that $na>b$. An element $a\in A$ is said to be non-negative if $a\geq 0$.
@@
Fix any $z\in H$ and $\bar{c}\in T\in\setminus\{0\}$. Write $c=(\bar{c},e_{G})$, and for every $n\in\omega$, let $nc=(n\bar{c},e_{G})$.
@@ 5. Comments and questions
Due to the restrictions imposed by [8, Corollary 12, Corollary 14] and Proposition 4.1, a natural candidate for the question above is a subsemigroup of $\mathbb{Q}^{(\mathfrak{c})}$. We note that Bellini, Rodrigues and Tomita [2] showed that for every free ultrafilter $p$, this group admits a $p$-compact group topology, unlike the free Abelian group which cannot have its $\omega$-th power countably compact [8].
@@ 6. Acknowledgements
The first author has received financial support from CAPES (Brazil) as PhD student at the University of São Paulo under supervision of the second listed author.
The second author is a voluntary (unpaid) postdoctoral researcher at the University of São Paulo. He thanks the University for the hospitality. He thanks the financial support of his parents Alberto Rodrigues, Tänia de Oliveira and his wife Bruna Nagano.

### 07 Mar 2024 v2

@@ 1.2. Basic results and notation
A free ultrafilter $p$ is said to be selective (or Ramsey) if for every $f:[\omega]^{2}\rightarrow 2$ there exists $A\in p$ such that $f|_{[A]^{2}}$ is constant. It is well known that selective ultrafilters are exactly the minimal ultrafilter in the Rudin-Keisler order, which is the pre-order defined in the collection of all free ultrafilters defined by $p\leq_{\operatorname{RK}}q$ if there exists a function $f:\omega\to\omega$ such that $p=\{f^{-1}[A]:A\in q\}$. Two selective ultrafilters are said to be incomparable if they are incomparable with respect to $\leq_{\operatorname{RK}}$. The existence of selective ultrafilters is known to be independent from ZFC.
@@ 4. $\mathbb{H}$ has no semigroup topology which makes its $\omega$-th power countably compact
We prove something more general. For that purpose, we say that an Archimedian semigroup with identity is a $4$-uple $(T,+,0,\leq)$ such that $(T,+,0)$ is an Abelian semigroup with identity 00, $\leq$ is a total order such that for all $a,b,c,d\in A$, if $a\leq b$ and $c\leq d$, then $a+c\leq b+d$, and for all $a>0$ and $b\in T$, there exists $n\in\mathbb{N}$ such that $na>b$. An element $a\in T$ is said to be non-negative if $a\geq 0$.
@@
Fix any $z\in H$ and $\bar{c}\in T\setminus\{0\}$. Write $c=(\bar{c},e_{G})$, and for every $n\in\omega$, let $nc=(n\bar{c},e_{G})$.
@@ 5. Comments and questions
Due to the restrictions imposed by [8, Corollary 12, Corollary 14] and Proposition 4.1, a natural candidate for the question above is a subsemigroup of $\mathbb{Q}^{(\mathfrak{c})}$. We note that Bellini, Rodrigues and Tomita [2] showed that for every selective ultrafilter $p$, this group admits a $p$-compact group topology, unlike the free Abelian group which cannot have its $\omega$-th power countably compact [8].
@@ 6. Acknowledgements
The first author has received financial support from CAPES (Brazil) as PhD student at the University of São Paulo under supervision of the third listed author.
The second author is a voluntary (unpaid) postdoctoral researcher at the University of São Paulo. He thanks the University for the hospitality. He thanks the financial support of his parents Alberto Rodrigues, Tânia de Oliveira and his wife Bruna Nagano.